OpenStudy (jojokiw3):

Can someone explain to me why the answer in this trigonometry question is what it is instead of something else?

OpenStudy (jojokiw3):

\[\sec (-\frac{ 19\pi }{ 6 })+\cos \frac{ 7\pi }{ 2 } - \sin \frac{ 10\pi }{ 3 } = \frac{ \sqrt{3} }{ 6 }\] instead of \[= -\frac{ 7\sqrt{3} }{ 6 }\]?

OpenStudy (jojokiw3):

\[-\frac{ \sqrt{3} }{ 6 }*\]

jimthompson5910 (jim_thompson5910):

can you provide what you got for each little piece? like what is the exact value of sec(-19pi/6) ? what is the exact value of cos(7pi/2) ? etc etc

OpenStudy (jojokiw3):

Wait. Lemme think. It's that way because -sin 10pi/3 falls into the 3rd quadrant which makes it a double negative instead of one negative?

OpenStudy (jojokiw3):

Because I got. -2/sqrt3 + 0 - sqrt3/2 whereas it should be -2/sqrt3 + 0 - (-sqrt3/2)?

jimthompson5910 (jim_thompson5910):

sin(10pi/3) = -sqrt(3)/2 so -sin(10pi/3) = -(-sqrt(3)/2) = sqrt(3)/2 you are correct

OpenStudy (jojokiw3):

Ahh. I forgot to put the quadrants in consideration. Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome